# Federated Learning and Differential Privacy Techniques on Multi-hospital Population-scale Electrocardiogram Data

### v1

@@ 1. Introduction
Electrocardiograms (ECGs), capturing the heart’s electrical activity, offer critical insights into cardiac function, making them an indispensable diagnostic and monitoring tool for healthcare professionals in cardiology. As this test is non-invasive, painless, and quick, it is a popular diagnostic tool in clinical settings, and is used to diagnose a range of cardiac conditions, including arrhythmias, conduction disorders, and other anomalies (7897225, ; Serhani2020-zu, ; 8946417, ; Liu2021-qa, ; Somani2021-gf, ). Recent studies show that ECGs can to identify and diagnose non-cardiac medical conditions, including but not limited to diabetes and Alzheimer’s disease (Sun2022-nx, ; Lou2023-sp, ; Ahn2022-hm, ). The manual analysis of ECG signals is a meticulous and error-prone task, even for seasoned specialists (Rafie2021-cu, ). Consequently, developing accurate ECG classification models is critical for the prompt and precise detection of cardiac conditions, which is vital for effective treatment and management. This necessity has spurred the adoption of deep learning techniques to leverage the full potential of ECG data. By automatically learning hierarchical representations from raw ECG signals, deep neural networks can uncover intricate features that might elude traditional handcrafted algorithms. Recent years have witnessed notable advancements in Machine Learning models for diagnosing diseases through ECG data. Employing various learners, including Convolutional Neural Networks (CNNs) (OShea2015-bz, ), Recurrent Neural Networks (RNNs) and their variants (Sherstinsky2020-jb, ), these learned models, particularly CNNs and their derivatives, excel in identifying spatial patterns, waveforms, and morphological features linked to specific medical conditions. Applying these deep learning techniques improves ECG classification accuracy, advancing cardiac diagnostics.
However, building such models requires access to substantial and diverse datasets that represent different patient populations and medical contexts. Data acquisition from multiple healthcare institutions is crucial to achieving generalizable models across various clinical and demographic populations. This maximizes model performance for the collaborating hospitals and solves the problems that may arise from having less training data. It provides a well-performing local model for a hospital for data from various clinical and demographic groups (Zhu2021-np, ; Prayitno2021-qk, ; Sadilek2021-mm, ). However, the acquisition of ECG data from many hospitals introduces intricate challenges and privacy considerations. The autonomous nature of hospitals leads to many different data formats, recording systems, and ECG acquisition protocols. Additionally, certain healthcare facilities may cater exclusively to specialized clinical cohorts or serve geographically confined populations. For instance, a predictive model optimized for a large urban hospital’s patient data, might perform less effectively when applied to the data from a small rural clinic, where the patient population and health issues differ significantly. Consequently, a model derived from one hospital’s dataset might not work for another hospital’s dataset due to the presence of covariate shifts across datasets from disparate institutions. A covariate shift between two hospitals’ data denotes a dissimilarity in patient characteristics, such as demographic features or disease prevalence, between the training dataset utilized for model development and the dataset encountered during deployment, potentially resulting in a decline in model performance.
Addressing privacy is an important concern, as hospitals are bound by a mandate to safeguard patient confidentiality and uphold ethical principles (aha, ; AMA, ; hipaa, ). Recent research has additionally demonstrated that ECGs can be leveraged for biometric authentication, showcasing their capacity to discern and verify individuals (Bras2018-ln, ; Melzi2023-cb, ). Furthermore, given the heightened sensitivity and confidentiality of medical data, specific individuals may be reluctant to provide their data to a central data collection (Ji2014-xq, ; Van_Zoonen2016-vo, ). The indiscriminate sharing of raw ECG data may inadvertently compromise sensitive information, eroding trust. Thus, the establishment of collaborative, privacy-centric data-sharing frameworks assumes paramount significance, with innovative methodologies such as Federated Learning (FL) and Differential Privacy (DP) emerging as promising solutions. This paper explores the application of FL and DP techniques on population-scale ECG data to develop accurate and robust ECG classification models while preserving data privacy and security.
@@
Sections 1.1 and 1.2 of this paper provides an overview of FL and DP, along with outlining the study’s objectives. Section 2 then delves into related literature, exploring FL and DP applications in different healthcare contexts, especially for ECG data. We critically assess the limitations of prior work and highlight our contributions. Section 3 outlines the methods employed, discussing the prediction task, patient characteristics, learning algorithm, and the evaluation scheme. Section 4 presents the results derived from various experiments conducted related to both FL and DP. Section 5 interprets our work and summarizes the entire paper.
@@ 1.1. Federated Learning
Federated Learning (FL) is a pioneering paradigm designed to address the multifaceted challenges associated with learning from many data sources in a way that preserves privacy (pmlr-v54-mcmahan17a, ). This new method allows one model to be trained together using data from different sources, like hospitals, without sharing private information. Each hospital helps improve that single model through minor updates while keeping sensitive details private. FL ensures that each hospital’s data stays within its boundaries, preserving data privacy and security. Figure 1 depicts the FL configuration employed in our multi-hospital setup.
@@
A woman and a girl in white dresses sit in an open car.
FL is a multi-step process in which each hospital initially preprocesses its local ECG data, a crucial step to ensure data consistency and implement privacy-preserving techniques, such as data anonymization. Subsequently, each hospital learns its own localized ECG classification model, based on its preprocessed data. These individual models are then aggregated into a global model through techniques such as federated averaging, which appropriately weighs each hospital’s contribution. The global model is then transmitted back to the individual hospitals, where it undergoes fine-tuning in successive iterations, until it converges on a global ECG classification model. Through this approach, FL establishes itself as a solution for collaborative medical research across diverse healthcare institutions.
@@ 1.2. Differential Privacy
A woman and a girl in white dresses sit in an open car.
Differential privacy (DP), a fundamental concept in data privacy, has become a crucial framework for addressing the delicate balance between preserving data usefulness and safeguarding individual privacy. It constitutes a foundational construct in the context of data analysis, which frequently involves using sensitive personal information while adhering to stringent ethical and legal mandates safeguarding the privacy of individuals. Initially introduced by Dwork et al. in 2006 (Dwork2006-al, ), this framework has since gained significant prominence as it offers a robust mathematical guarantee of privacy protection within data analysis processes. By quantifying the extent to which the inclusion or exclusion of a single individual’s data point influences the outcomes of data analyses, DP approaches to ensure that the privacy of any individual remains effectively preserved. In doing so, it furnishes a rigorous means of protecting individual privacy and inspires confidence in data sharing and analysis, thus facilitating collaborative medical research and advancing healthcare outcomes. Figure 2 illustrates the DP setup in the context of machine learning. In this scenario, we can train 2 models using learner A: model M on database D and M’ on database D’, which is the dataset formed by adding Mr. Smith’s data to dataset D. Subsequently, we pose the same query Q to both of these trained models. If the resulting answers, B and B’, are indistinguishable, then no private information about Mr. Smith can be inferred.
@@ 1.3. Aim of Study
This study investigates the impact of incorporating FL and DP techniques in the analysis of ECG data obtained from 7 distinct hospitals located in Edmonton, Alberta, Canada. Our prediction task is to develop an ECG classification model that can use a patient’s ECG to determine if that patient has zero or more of several prevalent cardiovascular and metabolic diseases. The overarching goal is to investigate whether adopting a federated approach with (versus without) differential privacy can enhance data privacy and security without significantly compromising the performance of pre-existing global models.
@@ 2. Related Work
Various studies have explored the implementation of FL in healthcare applications. One such study by Raza et al. (Raza2022-an, ) employed an FL approach to categorize ECGs into 5 heartbeat categories on the MIT-BIH Dataset (Moody2001-ib, ) focusing on model explainability. However, it uses data from a single hospital, undermining the fundamental justification for employing FL in this scenario. In contrast, our research overcomes this limitation by incorporating data aggregated from multiple hospitals. In a customized FL approach, Tang et al. (Tang2021-mq, ) demonstrates enhanced performance in their FL model, personalized for each local node, compared to a model trained using the FedAvg algorithm. However, their focus — classifying ECGs into distinct heartbeat types —is designed to be relevant to the diseases under consideration. Additionally, their study would benefit from a comparative analysis across demographic groups and an inter-site performance evaluation. Our study used the same FedAvg algorithm in an inter-site analysis framework inspired by the work conducted by Goto et al. (Goto2022-ut, ). They developed FL models for hypertrophic cardiomyopathy classification by integrating ECGs and echocardiograms from 4 different hospitals, thereby setting a benchmark for assessing the utility of ECGs. We differ from the HCM-focused study (Goto2022-ut, ) as we leverage ECG data to comprehensively diagnose various cardiovascular and metabolic diseases in patients — more details in Sections 3.1 and 3.2. Both Lin et al. (9797945, ) and Meqdad et al. (Meqdad2023-vh, ) propose novel FL approaches for 5-class heartbeat classification using ECGs, specifically beneficial for non-IID smart device data, with validation on the MIT-BIH dataset. The non-IID nature of the data in this context arises because the data from each node (patient or device) originates from the same patient, leading to a lack of independence and thus categorizing it as non-IID. In contrast to Lin et al. (9797945, ) and Meqdad et al. (Meqdad2023-vh, ), Baumgartner et al. (Baumgartner2023-fn, ) learns a multi-label classification model that can detect 13 SNOMED codes using ECGs from 5 data sources (moody, ), emphasizing metric-focused evaluations while using FL to preserve privacy for multi-site data. Ying et al. (Ying2023-gi, ) introduces a pioneering preprocessing method for ECG data by converting ECGs to images and employs FL for smart devices, incorporating a semi-supervised approach with pseudo-labeling. Nevertheless, they conducted an analysis using the MIT-BIH dataset, known for its non-IID nature and utility for patient-level data nodes. Joshi, Pal et al. (Joshi2022-zh, ) provides a comprehensive review of prevailing FL approaches in healthcare until 2022. Finally, Dolo et al. (10017908, ) proposes for the Differentially Private Stochastic Gradient Descent applied to the Federated Averaging (DPSGDFedAvg) algorithm, specifically for diabetes prediction using the Pima Indian dataset. We adopt DPSGDFedAvg for our experiments using DP in Section 4.2.
@@
This study adds to the existing body of literature by introducing several advancements, thus distinguishing itself from prior research. An essential point of distinction lies in the fact that numerous papers utilize ECG data from the MIT-BIH dataset sourced from a single hospital. In these cases, FL is implemented on an individual patient level, treating them as nodes, and the data is non-IID. In contrast, our approach uses hospitals as nodes in our FL model, mirroring a more realistic deployment scenario. Our model utilizes data from 7 hospitals in Edmonton, Alberta, Canada, where each hospital contains data from numerous patients. The linkage of our ECG dataset to provincial administrative health records provides multiple diagnostic labels based on ICD-10 coding systems, meaning our learned models can predict many more labels than others that incorporate private data from multiple sources (Baumgartner2023-fn, ; Tang2021-mq, ). Our study ensures a substantial number of instances for each disease type, with a minimum of 15,000 instances for the least prevalent disease, facilitating ample training samples. Additionally, a distinctive characteristic of our machine learning approach is that our learned model will simultaneously predict the diagnosis of multiple diseases for a given ECG instance. We employ the FedAvg algorithm for FL, consistent with the existing literature, considering its established efficacy. Furthermore, we introduce an additional layer of data privacy by incorporating the Differential Privacy Stochastic Gradient Descent (DP-SGD) algorithm (Abadi2016-fu, ) during FL training. The inclusion of DP-SGD is a novel aspect absent in previous related works on similar prediction tasks and aligns with contemporary privacy norms and legal requirements. Additionally, our research conducts intersite analysis, addressing a notable gap in some previous literature. This analysis underscores the motivation behind employing FL, particularly for hospitals with limited ECG data, thus ensuring both efficacy and data privacy simultaneously, a crucial aspect not explored in earlier studies. Furthermore, we aim to demonstrate that the effective implementation of the FL approach is influenced by factors such as the distribution of disease prevalence among different hospitals, the age-sex distribution, and the number of ECGs available in these hospitals. In addition, this study seeks to showcase the trade-off between model performance and data security by employing DP techniques in model training.
@@ 3.1. Data Description and Patient Characteristics
This work used a dataset comprising 2,015,808 ECG records from 260,065 patients, collected between February 2007 and April 2020, containing 14 hospitals in Edmonton, Alberta, Canada. We have implemented identical ECG data preprocessing methods as outlined in the Analysis Cohort subsection of the Methods section in Sun et al. (Sun2023-lp, ). An ECG record was labeled with ICD-10 codes of the healthcare episode if its acquisition date fell within the timeframe of the episode. After excluding ECGs that could not be linked to any episode, those of patients below 18 years of age, and ECGs with poor signal quality, our analysis cohort comprised 1,603,109 ECGs originating from 748,773 episodes involving 244,077 patients (Sun2023-lp, ).
@@
Table 1 provides the distribution of ECGs across various hospitals along with different experimental splits. We have set a training dataset quota of 12,500 ECGs per hospital for model development–meaning our training set, used to train our Federated Learning-based model, included just 7 hospitals. The final dataset consists of 1,565,849 ECGs, encompassing 243,128 patients. This study employs standard 12-lead ECG tracings derived from the Philips IntelliSpace ECG system, consisting of voltage-time series, sampled at a rate of 500 Hz over a duration of 10 seconds for each of the 12 leads, resulting in a total of 500 × 10 × 12 voltage measurements per ECG. General patient characteristics have been described by Sun et al. (Sun2023-lp, ), and hospital-wise distributions are presented in Table 2 and Table 3.
Table 2. Patient Characteristics
@@
Table 3. Prevalences of classification labels and comorbidities
Hospital 1 2 3 4 5 6 7
@@ 3.2. Prediction Task
Our dataset, sourced from Alberta Health Services (AHS), contains administrative electronic health records (EHRs) and ECGs. Based on linkage of ECGs to administrative health records, we labeled each ECG with the set of ICD-10 diagnosis codes assigned to this patient. The objective of our research is a multi-label classification task that uses both ECG waveforms and demographic data to predict the probabilities of 10 specific diseases, as defined by their respective ICD-10 codes (see Table 6). We chose these classification labels because of their significant clinical relevance to cardiovascular and metabolic diseases and their widespread prevalence across diverse demographic cohorts, as elucidated by the research findings presented by Sun et al. (Sun2023-lp, ). Figure 3 provides an example of the predictions generated by our model. The model predicts each disease label independently, assigning either 0 (negative) or 1 (positive) for a given ECG tracing as input. Our dataset aggregates information from multiple hospitals, so our approach involves developing a multi-label model within a federated learning framework. We compare the standard and federated learning models to discern the efficacy of federated learning in this context. Finally, we will compare models trained and tested independently within and across each participating hospital, evaluating their performance.

### v2

@@ 1. Introduction
Electrocardiograms (ECGs), offer critical insights into cardiac function by capturing the heart’s electrical activity, making them an indispensable tool in healthcare. As this test is non-invasive, painless, and quick, it is used to diagnose a range of cardiac conditions, including arrhythmias, conduction disorders, and other anomalies (Liu2021-qa, ; Serhani2020-zu, ). Recent studies show that ECGs can also identify non-cardiac medical conditions, including diabetes and Alzheimer’s disease (Ahn2022-hm, ; Sun2022-nx, ). Unfortunately, the manual analysis of ECG signals is a meticulous and error-prone task, even for seasoned specialists (Rafie2021-cu, ). Consequently, developing accurate ECG classification models is critical for the prompt and precise detection of medical conditions, which is vital for effective treatment and management. This necessity has spurred the adoption of deep learning techniques to leverage the full potential of ECG data. By automatically learning hierarchical representations from raw ECG signals, deep neural networks can uncover intricate features that might elude traditional handcrafted algorithms. However, building such models requires access to substantial and diverse datasets that represent different patient populations and medical contexts to achieve generalizable models across various clinical and demographic populations. Moreover, the acquisition of ECG data from many hospitals introduces challenges of privacy considerations. The autonomous nature of hospitals leads to many different data formats, recording systems, and ECG acquisition protocols. Additionally, certain healthcare facilities may cater exclusively to specialized clinical cohorts or serve geographically confined populations. A covariate shift between two hospitals’ data denotes a dissimilarity in patient characteristics, such as demographic features or disease prevalence, between the training dataset utilized for model development and the dataset encountered during deployment, potentially resulting in poor model performance. Concurrently, addressing privacy is important, as hospitals are bound by a mandate to safeguard patient confidentiality and uphold ethical principles (aha, ; AMA, ; hipaa, ). Recent research has additionally demonstrated that ECGs can be leveraged for biometric authentication, showcasing their capacity to discern and verify individuals (Bras2018-ln, ; Melzi2023-cb, ). Furthermore, given the heightened sensitivity and confidentiality of medical data, certain individuals may be reluctant to provide their data to a central data collection (Ji2014-xq, ; Van_Zoonen2016-vo, ). Thus, the establishment of collaborative, privacy-centric data-sharing frameworks assume paramount significance, with innovative methodologies such as Federated Learning (FL) and Differential Privacy (DP) emerging as promising solutions. It is essential to recognize that the FL method confers significant advantages in bolstering data security and guaranteeing privacy, as delineated by Yin et al. (Yin2022-st, ). This paper explores the application of FL and DP techniques on population-scale ECG data to develop accurate ECG classification models while preserving data privacy and security.
@@
Subsections 1.1 and 1.2 provide an overview of FL and DP, and outline the study’s objectives. Section 2 then delves into related literature, exploring FL and DP applications in different healthcare contexts, especially for ECG data –to critically assess the limitations of prior work and highlight our contributions. Section 3 outlines the methods we will employ, discusses the prediction task, patient characteristics, learning algorithm, and the evaluation scheme. Section 4 presents the results derived from various experiments conducted related to both FL and DP. Section 5 interprets our work and summarizes the entire paper.
@@ 1.1. Federated Learning
Federated Learning (FL) is a pioneering paradigm designed to address the challenges associated with learning from many data sources in a way that preserves privacy (pmlr-v54-mcmahan17a, ). This method allows one model to be trained using data from different sources, like hospitals, without sharing private information. Each hospital helps improve that single model through minor updates while keeping sensitive details private. FL ensures that each hospital’s data stays within its boundaries, preserving data privacy and security. Figure 1 depicts the FL configuration employed in our multi-hospital setup, showing FL is a multi-step process in which each hospital initially preprocesses its local ECG data, a crucial step to ensure data consistency and implement privacy-preserving techniques. Subsequently, each hospital learns its own localized ECG classification model, based on its preprocessed data. These individual models are then aggregated into a global model through techniques such as federated averaging, which appropriately weighs each hospital’s contribution (pmlr-v54-mcmahan17a, ). The global model is then transmitted back to the individual hospitals, where it undergoes fine-tuning in successive iterations; these “weights” are then returned to the central, until it converges on a single global ECG classification model. Through this approach, FL establishes itself as a solution for collaborative medical research across diverse healthcare institutions.
@@ 1.2. Differential Privacy
Differential privacy (DP), a fundamental concept in data privacy, has become a crucial framework for addressing the delicate balance between preserving data usefulness and safeguarding individual privacy. Initially introduced by Dwork et al. in 2006 (Dwork2006-al, ), this framework has since gained prominence as it offers a mathematical guarantee of privacy protection within data analysis processes. By quantifying the extent to which the inclusion or exclusion of a single individual’s data point influences the outcomes of data analyses, DP ensures that the privacy of any individual remains effectively preserved, inspiring confidence in data sharing, which facilitates collaborative medical research for advancing healthcare outcomes. Figure 2 illustrates the DP setup in the context of machine learning. In this scenario, we can train 2 models using learner A: model M on database D and M’ on database D’, which is the dataset formed by adding Mr. Smith’s data to dataset D. Subsequently, we pose the same query Q to both of these trained models. If the resulting answers, B and B’, are essentially indistinguishable, then no private information about Mr. Smith can be inferred.
@@ 1.3. Aim of Study
Our study investigates the impact of incorporating FL and DP techniques in the analysis of ECG data obtained from 7 distinct hospitals in Alberta, Canada. Our prediction task is to develop an ECG classification model that can use a patient’s ECG to determine if that patient has zero or more of several prevalent cardiovascular and metabolic diseases. The overarching goal is to investigate whether adopting a federated approach with (versus without) differential privacy can enhance data privacy and security without significantly compromising the performance of pre-existing global models.
@@ 2. Related Work
Various studies have explored the implementation of FL in healthcare applications (Joshi2022-zh, ). Raza et al. (Raza2022-an, ) employed an FL approach on the MIT-BIH Dataset (Moody2001-ib, ), to learn a model to categorize ECGs into 5 heartbeat categories focusing on model explainability. However, it uses data from a single hospital, undermining the fundamental justification for employing FL in this scenario. In contrast, our research matches the FL objective by incorporating data aggregated from multiple hospitals. In a customized FL approach, Tang et al. (Tang2021-mq, ) demonstrates enhanced performance in their FL model, personalized for each local node, compared to a model trained using the FedAvg algorithm. Their study would benefit from a comparative analysis across demographic groups and an inter-site performance evaluation. Our study used the same FedAvg algorithm in an inter-site analysis framework inspired by Goto et al. (Goto2022-ut, ), who developed FL models for hypertrophic cardiomyopathy classification by integrating ECGs and echocardiograms from 4 different hospitals, thereby setting a benchmark for assessing the utility of ECGs. We differ from the HCM-focused study (Goto2022-ut, ) as we leverage ECG data to comprehensively diagnose various cardiovascular and metabolic diseases in patients — more details in Subsections 3.1 and 3.2. Both Lin et al. (9797945, ) and Meqdad et al. (Meqdad2023-vh, ) propose novel FL approaches for 5-class heartbeat classification using ECGs, specifically beneficial for non-independent and identically distributed (non-IID) smart device data, with validation on the MIT-BIH dataset. The non-IID nature of the data in this context arises because the data from each node (patient or device) originates from the same patient, leading to a lack of independence and thus categorizing it as non-IID. In contrast, Baumgartner et al. (Baumgartner2023-fn, ) learns a multi-label classification model that can detect 13 SNOMED codes using ECGs from 5 data sources (moody, ), emphasizing metric-focused evaluations while using FL to preserve privacy for multi-site data. Ying et al. (Ying2023-gi, ) introduces a novel preprocessing method for ECG data by converting ECGs to images and employs FL for smart devices, incorporating a semi-supervised approach with pseudo-labeling. Nevertheless, they conducted an analysis using the MIT-BIH dataset, known for its non-IID nature and utility for patient-level data nodes. Finally, Dolo et al. (10017908, ) proposes for the Differentially Private Stochastic Gradient Descent applied to the Federated Averaging (DPSGDFedAvg) algorithm, specifically for diabetes prediction. We adopt DPSGDFedAvg for our experiments in Section 4.3.
@@
Our study adds to the existing body of literature by introducing several advancements. An essential difference is that numerous papers utilize ECG data from the MIT-BIH dataset sourced from a single hospital. In these cases, FL is implemented on an individual patient level, treating them as nodes, and the data is non-IID. In contrast, our approach uses hospitals as nodes in our FL model, mirroring a more realistic deployment scenario. Our model utilizes data from 7 hospitals in Alberta, Canada, where each hospital contains data from numerous patients. The linkage of our ECG dataset to provincial administrative health records provides multiple diagnostic labels based on ICD-10 coding systems and our learned model can simultaneously predict the diagnosis of multiple diseases for a given ECG instance (Baumgartner2023-fn, ; Tang2021-mq, ). Our study focuses on just diseases that include at least 15,000 instances, facilitating ample training samples. We employ the FedAvg algorithm (pmlr-v54-mcmahan17a, ) for FL, consistent with the existing literature, considering its established efficacy. Furthermore, we introduce an additional layer of data privacy by incorporating the Differential Privacy Stochastic Gradient Descent (DP-SGD) algorithm (Abadi2016-fu, ) during FL training. The inclusion of DP-SGD differs from previous related works on similar prediction tasks and aligns with contemporary privacy norms and legal requirements. Additionally, our research conducts inter-site analysis, addressing a notable gap in some previous literature. This analysis underscores the motivation behind employing FL, particularly for hospitals with limited ECG data, thus ensuring both efficacy and data privacy simultaneously, a crucial aspect not explored in earlier studies. In addition, this study seeks to showcase the trade-off between model performance and data security by employing DP techniques in model training.
@@ 3.1. Data Description and Patient Characteristics
This work began with a dataset containing 2,015,808 ECG records from 260,065 patients, collected between February 2007 and April 2020, from 14 hospitals in Edmonton, Alberta, Canada. We applied the ECG data preprocessing methods outlined in the Analysis Cohort subsection of the Methods section in Sun et al. (Sun2023-lp, ). An ECG record was labeled with ICD-10 codes of the healthcare episode if its acquisition date fell within the timeframe of the episode. We excluded ECGs that (1) could not be linked to any episode, (2) were from patients below 18 years of age, or were of poor signal quality, leaving an analysis cohort of 1,603,109 ECGs originating from 748,773 episodes involving 244,077 patients (Sun2023-lp, ). In this study, we included 7 hospitals (also referred to as ‘sites’), each of which provided a minimum of 12,500 ECGs. Table 1 provides the distribution of ECGs across various hospitals along with different experimental splits. The final dataset consists of 1,565,849 ECGs, from 243,128 patients across 7 hospitals. Here, each ECG is a standard 12-lead ECG tracings derived from the Philips IntelliSpace ECG system, consisting of voltage-time series, sampled at a rate of 500 Hz over a duration of 10 seconds for each of the 12 leads, resulting in a total of 500 × 10 × 12 voltage measurements per ECG. General patient characteristics have been described by Sun et al. (Sun2023-lp, ), and hospital-wise distributions are presented in Table 2.
@@
Table 2. Patient Characteristics and Prevalence of classification labels and comorbidities
@@ 3.2. Prediction Task
Our dataset, sourced from Alberta Health Services (AHS), contains administrative electronic health records (EHRs) and ECGs. Based on linkage of ECGs to administrative health records, we labeled each ECG with the set of ICD-10 diagnosis codes assigned to this patient. The objective of our research is a multi-label classification task that uses both ECG waveforms and demographic data to predict the probabilities of 10 specific diseases, as defined by their respective ICD-10 codes (see Appendix: Table 5). We chose these classification labels because of their significant clinical relevance to cardiovascular and metabolic diseases and their predictability using Deep Learning (DL) models, as elucidated by the research findings presented by Sun et al. (Sun2023-lp, ). Figure 3 provides a schematic of the performance system generated by our model. The model predicts each disease label independently, assigning either 0 (negative) or 1 (positive) based on the prediction probability for a given ECG tracing as input. Our dataset aggregates information from multiple hospitals, so our approach involves developing a multi-label model within a federated learning framework. We compare the pooled and federated learning models to discern the efficacy of federated learning in this context. Finally, we will compare models trained and tested independently within and across each participating hospital, evaluating their performance.